# MOA-2010-BLG-328: Keck and HST Expose the Limits of Occams Razor in Microlensing

### v1

@@ 5.1 Investigating the degenerecy
In Figure 6 we see two possible solutions with either an M-dwarf host of $\sim 0.2\ M_{\odot}$ at $\sim 2$ kpc or a higher mass one with $\sim 0.5\ M_{\odot}$ at $\sim 4$ kpc. This degeneracy is characteristic for M-dwarfs in the disk where different host masses can be challenging to distinguish since an early M-dwarf further away can appear to be a late M-dwarf that is closer.
@@
Due to the interplay between all the higher order effects of microlensing parallax signals, xallarap, magnification of the second source and planetary orbital motion ambiguity is created. When we examine the corner plot in Figure 7 of the higher order effects, we clearly see that the microlensing parallax is not constrained, even with the direct measurement of the source-lens relative proper motion. We investigate these $\pi_{\rm E}$ values by making a diagonal cut in the ($\pi_{\rm E,E}$, $\pi_{\rm E,N}$) space, at $\pi_{\rm E,E}=\pi_{\rm E,N}-0.33$. This can be seen on Figure 6 where orange represents the “larger” values where $\pi_{\rm E}>\pi_{\rm E,N}-\pi_{\rm E,E}-0.33$, and green represents the “smaller” values where $\pi_{\rm E}<\pi_{\rm E,N}-\pi_{\rm E,E}-0.33$.
@@
This is also shown in Figure 8 which shows the mass-distance and mass-luminosity constraints from the high resolution follow-up data and described in Equations 3, 4 and 7. In this type of Figure, the intersection of different constraints results in the mass and distance of the lens star. In Panel A of this Figure, however, we show that without the addition of a microlensing parallax measurement, there is a continuous degeneracy where the mass-distance constraint due to $\theta_{E}$ and that of due to the lens $K$ magnitude, do not intersect in any one place. The addition of a parallax measurement should break this degeneracy, however, as we saw in Figure 7 there are two possible solutions for the microlensing parallax - which leads to two possible solutions for the lens mass and distance. We plot these separately in Panels B and C of Figure 8
We also investigate the expected $V$ and $I$ magnitude of the lens associated with the large and small $\pi_{\rm E}$ values from the MCMC probability distribution. We find the following $2\sigma$ ranges ($26.2825.68<V_{\rm L}<$ 26.28), ($22.5021.96<I_{\rm I}<$ 22.50) and ($25.5923.50<V_{\rm L}<$ 25.59), ($22.1820.81<I_{\rm I}<$ 22.18) for the large and small $\pi_{\rm E}$ values, respectively. This means a V-band gap of 0.084 and an I-band gap of -0.221, which means this degeneracy can be broken with additional data in shorter-wavelengths, with telescopes such as HST or JWST.
Another contributing factor is the lack of coverage at the beginning of the planetary deviation, which can be seen in the middle panel of Figure 5. This has resulted in large uncertainties on the host-planet mass ratio, which explains the large error bars on the planet mass.
Figure 6: Bayesian probability distributions of the planetary system, including host mass and distance, separation and planetary companion mass. Orange represents $\pi_{\rm E}>\pi_{\rm E,N}-\pi_{\rm E,E}-0.33$ values, and green respresents $\pi_{\rm E}<\pi_{\rm E,N}-\pi_{\rm E,E}-0.33$ values. The central 68.3% of the distributions are shaded in darker hues (of either orange or green) and the remaining central 95.4% of the distributions are shaded in lighter hues. The vertical black line marks the median of the probability distribution of the respective parameters for each $\pi_{\rm E}$ above and below the cut.
Figure 7: Corner plot from the MCMC probability distribution showing the parameters due to the microlenisng parallax, orbital motion, xallarp and magnification of the secondary source.
@@ 6 Discussion and Conclusion
Therefore, in this paper we remodel the light curve using all these higher order effects; microlensing parallax, orbital motion, xallarap, and magnification of the second source. In the initial study by Furusawa et al. (2013) they explored higher order effects on the light curve, but did not present a model that combined all of them. Instead, they investigated an orbital motion plus parallax model, and a xallarap model separately, finding these two to be degenerate. We also include constraints from the high resolution follow-up data from Keck and HST. From Keck we use the measured magnitude and direction of the relative source-lens proper motion as a constraint, and from HST we use the total (source+lens) measured $V$ and $I$ magnitudes as constraints. This new model was able to both match the positions and magnitudes of the lens and source stars, as seen in the Keck and HST data. The higher order constraints in this case play an important role when modeling the light curve. Modeling of planetary systems is in general simple if constraints are applied, and typically acts to improve the efficiency of the MCMC by excluding models that are inconsisted with the high resolution follow-up observations. However, in cases like MB10328 where parallax, orbital motion and xallarap are applied, without constraints from high resolution observations it is possible that the modeling code would be unable to find the correct solution. The likelihood of a model to fail is increased when the higher order effect parameters are not constrained to physically reasonable values with a prior distribution.
@@
MB10328 is also part of the Suzuki et al. (2016) statistical sample, and these results will contribute to the determination of exoplanet host star masses. The Suzuki et al. (2016) sample will help us understand the dependence of planet occurance rate on the mass and distance of the host star. The analysis of other events in this sample have also highlighted the importance of including higher order effects on the light curve. Our high angular resolution follow-up data (from Keck and HST) indicate that the modeling of these events had to be redone in order to determine the physical parameters of the system e.g. MOA-2007-BLG-192 (Bennett et al., 2008; Terry et al., 2024) and OGLE-2012-BLG-0563 (Fukui et al. (2015), Bennet et al. submitted to AJ, Bhattacharya et al. submitted to AJ).
@@
The inclusion of higher order effects like microlensing parallax in light curve modeling is of particular importance. Every microlensing event, observed in a heliocentric reference frame will have microlensing parallax present, even if the effect is not measurable on the light curve. Ignoring these effects could lead to errors in other light curve parameters. This poses challenges in determining the physical parameters of microlensing systems, which complicates conducting statistical analyses of exoplanetary system properties from microlensing surveys, such as the future Roman Galactic Exoplanet Survey (RGES) (Bennett et al., 2018; Penny et al., 2019; Johnson et al., 2020). In this work we show the importance of including higher order effects on the light curve and especially the use of image constrained modeling with high resolution follow-up data. This method will play an important role for the RGES to measure the masses and distances of all the exoplanetary systems it detects.

### v2

@@ 5.1 Investigating the degenerecy
In Figure 9 we see two possible solutions with either an M-dwarf host of $\sim 0.2\ M_{\odot}$ at $\sim 2$ kpc or a higher mass one with $\sim 0.5\ M_{\odot}$ at $\sim 4$ kpc. This degeneracy is characteristic for M-dwarfs in the disk where different host masses can be challenging to distinguish since an early M-dwarf further away can appear to be a late M-dwarf that is closer. The planetary companion’s mass is also affected. If the host is a late M-dwarf at a closer distance it is likely to host a sub-Neptune mass planet, if the host is a early M-dwarf further away, it likely hosts a Gas Giant with a mass $\sim 2\ M_{\rm Saturn}$. The projected separation for each solution can be calculated with $a_{\perp}=sD_{\rm L}\theta_{\rm E}$. For both solutions, the planetary companion is likely to be beyond the snow line.
@@
Due to the interplay between all the higher order effects of microlensing parallax signals, xallarap, magnification of the second source and planetary orbital motion ambiguity is created. This can be seen on the corner plot in Figure 7 of the higher order effects. For the lens system’s physical parameters this degeneracy can also be seen in Figure 9, where orange represents the “larger” values ( $\pi_{\rm E,N}>\pi_{\rm E,E}-0.33$), and green represents the “smaller” values ($\pi_{\rm E,N}<\pi_{\rm E,E}-0.33$).
@@
This is also shown in Figure 10 which shows the mass-distance and mass-luminosity constraints from the high resolution follow-up data and described in Equations 4, 5 and 8. In this type of Figure, the intersection of different constraints results in the mass and distance of the lens star. In Panel A of this Figure, however, we show that without the addition of a microlensing parallax measurement, there is a continuous degeneracy where the mass-distance constraint due to $\theta_{E}$ and that of due to the lens $K$ magnitude, do not intersect in any one place. The addition of a parallax measurement should break this degeneracy, however, as we saw in Figure 7 there are two possible solutions for the microlensing parallax - which leads to two possible solutions for the lens mass and distance. We plot these separately in Panels B and C of Figure 10
We also investigate the expected $V$ and $I$ magnitude of the lens associated with the large and small $\pi_{\rm E}$ values from the MCMC probability distribution. We find the following $2\sigma$ ranges ($25.68<V_{\rm L}<26.28$), ($21.96<I_{\rm I}<22.50$) and ($23.50<V_{\rm L}<25.59$), ($20.81<I_{\rm I}<22.18$) for the large and small $\pi_{\rm E}$ values, respectively. This means a V-band gap of 0.084 and an I-band gap of -0.221, which means this degeneracy can be broken with additional data in shorter-wavelengths, with telescopes such as HST or JWST.
Another contributing factor to this degeneracy is the lack of coverage at the beginning of the planetary deviation between $\sim$5400 - 5406 days, which can be seen in the middle panel of Figure 6. In Figure 11 we see that there is a small difference between the two models at the entrance of the planetary anomaly. The light-curve parameter affected the most by this is the host-planet mass ratio $q$, which is constrained by the shape and strength of the planetary anomaly.
@@
Figure 11: A zoom of the entry of the planetary anomaly for the two best-fit models, small $\pi_{\rm E}$ in orange and large $\pi_{\rm E}$ in green, and the corresponding MOA and OGLE data depicted by a circle and diamond, respectively.
@@ 6 Discussion and Conclusion
Therefore, in this paper we remodel the light-curve using all these higher order effects; microlensing parallax, orbital motion, xallarap, and magnification of the second source. In the initial study by Furusawa et al. (2013) they explored higher order effects on the light-curve, but did not present a model that combined all of them. Instead, they investigated an orbital motion plus parallax model, and a xallarap model separately, finding these two to be degenerate. We also include constraints from the high resolution follow-up data from Keck and HST. From Keck we use the measured magnitude and direction of the relative source-lens proper motion as a constraint, and from HST we use the total (source+lens) measured $V$ and $I$ magnitudes as constraints. This new model was able to both match the positions and magnitudes of the lens and source stars, as seen in the Keck and HST data. The higher order constraints in this case play an important role when modeling the light-curve. Modeling of planetary systems is in general simple if constraints are applied, and typically acts to improve the efficiency of the MCMC by excluding models that are inconsisted with the high resolution follow-up observations. However, in cases like MB10328 where parallax, orbital motion and xallarap are applied, without constraints from high resolution observations it is possible that the modeling code would be unable to find the correct solution. The likelihood of a model to fail is increased when the higher order effect parameters are not constrained to physically reasonable values with a prior distribution.
@@
MB10328 is also part of the Suzuki et al. (2016) statistical sample, and these results will contribute to the determination of exoplanet host star masses. The Suzuki et al. (2016) sample will help us understand the dependence of planet occurance rate on the mass and distance of the host star. The analysis of other events in this sample have also highlighted the importance of including higher order effects on the light-curve. Our high angular resolution follow-up data (from Keck and HST) indicate that the modeling of these events had to be redone in order to determine the physical parameters of the system e.g. MOA-2007-BLG-192 (Bennett et al., 2008; Terry et al., 2024) and OGLE-2012-BLG-0563 (Fukui et al. (2015); Bennett et al. (2024), Bhattacharya et al., submitted to AJ).
@@
The inclusion of higher order effects like microlensing parallax in light-curve modeling is of particular importance. Every microlensing event, observed in a heliocentric reference frame will have microlensing parallax present, even if the effect is not measurable on the light-curve. Ignoring these effects could lead to errors in other light-curve parameters. This poses challenges in determining the physical parameters of microlensing systems, which complicates conducting statistical analyses of exoplanetary system properties from microlensing surveys, such as the future Roman Galactic Exoplanet Survey (RGES) (Bennett et al., 2018; Penny et al., 2019; Johnson et al., 2020). In this work we show the importance of including higher order effects on the light-curve and especially the use of image constrained modeling with high resolution follow-up data. This method will play an important role for the RGES to measure the masses and distances of all the exoplanetary systems it detects.